# Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebula

### v1

@@ Abstract
Among over 1000 known fast radio bursts (FRBs), only three sources – FRB 121102 (R1), FRB 190520 (R2) and FRB 201124 (R3) – have been linked to persistent radio sources (PRS). The observed quasi-steady emission is consistent with synchrotron radiation from a composite of magnetar wind nebula (MWN) and supernova (SN) ejecta. We compute the synchrotron flux by solving kinetic equations for energized electrons, considering electromagnetic cascades of electron-positron pairs interacting with nebular photons. For rotation-powered model, a young neutron star (NS) with age $t_{\rm age}\approx 20\,{\rm yr}$, dipolar magnetic field $B_{\rm dip}\approx(3-5)\times 10^{12}\,{\rm G}$ and spin period $P_{i}\approx 1.5-3\,{\rm ms}$ in an ultra-stripped SN progenitor can account for emissions from R1 and R2. In contrast, R3 requires $t_{\rm age}\approx 10\,{\rm yr}$, $B_{\rm dip}\approx 5.5\times 10^{13}\,{\rm G}$ and $P_{i}\approx 10\,{\rm ms}$ in a conventional core-collapse SN progenitor. For magnetar-flare-powered model, NS aged $t_{\rm age}\approx 25\,/40\,{\rm yr}$ in a USSN progenitor and $t_{\rm age}\approx 12.5\,{\rm yr}$ in a CCSN progenitor explains the observed flux for R1/R2 and R3, respectively. Finally, we constrain the minimum NS age $t_{\rm age,min}\sim 1-3\,{\rm yr}$ from the near-source plasma contribution to observed DM, and $t_{\rm age,min}\sim 6.5-10\,{\rm yr}$ based on the absence of radio signal attenuation.
@@
stars: magnetars, winds, outflows – supernovae: general – transients: fast radio bursts – radiation mechanisms: general
pagerange: Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebulaQuasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebulapubyear: 2024
@@ 1 Introduction
Fast radio bursts (FRBs) are energetic millisecond duration pulses of coherent emission, located at cosmological distances, whose physical origin is still debated after a decade since their discovery (Lorimer et al., 2007; Thornton et al., 2013; Spitler et al., 2014; Cordes & Chatterjee, 2019). Many theoretical models have been proposed to explain the nature of their progenitors, both for repeating and one-off events (see Platts et al. 2019, for a recent review). The detection of Galactic FRB 200428 and its association with SGR J1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), suggests that FRBs can originate from magnetars born from the core collapse of massive stars (see e.g., Murase et al. 2016; Kumar et al. 2017; Bhattacharya 2019; Kumar & Bošnjak 2020; Lu et al. 2022).
@@
In addition to FRBs, rapidly rotating pulsars and magnetars have been proposed as central engines of super-luminous supernovae (SNe), stripped-envelope SNe and rapidly-rising optical transients (Metzger et al., 2015; Kashiyama et al., 2016; Hotokezaka et al., 2017; Margalit & Metzger, 2018). Pulsar wind nebulae (PWNe) are known to be efficient particle accelerators and broadband non-thermal emission has been observed from the nebulae of Galactic PWNe (Tanaka & Takahara, 2010, 2013). Murase et al. (2016) proposed quasi-steady synchrotron emission as counterparts of both FRBs and pulsar/magnetar-driven SNe (including SLSNe), suggesting efficient conversion of rotation and/or magnetic energy to particle energies within the nascent MWN of a young NS (e.g., Gaensler & Slane 2006).
@@
Precise localisation of FRBs can provide meaningful insights on their sources, by identifying plausible multi-wavelength counterpart(s) and revealing information about the central engine and its surrounding environment (Michilli et al., 2018). However, until date, only three confirmed FRB-PRS associations have been made in the sample of localized FRBs; namely those of FRB 121102 (Chatterjee et al., 2017), FRB 190520 (Niu et al., 2022) and FRB 201124 (Bruni et al., 2023). PRSs are continuum radio sources that are significantly bright ($L_{\rm PRS}>10^{29}\,{\rm erg\,s^{-1}\,Hz^{-1}}$) and compact ($R_{\rm PRS}<1\,{\rm pc}$), for them to be related to ongoing star formation in their host galaxy (Nimmo et al., 2022; Dong et al., 2024). The polarisation level is found to be significantly different between the FRB 121102 bursts and its PRS emission (Michilli et al., 2018; Plavin et al., 2022), thereby ruling out the possibility that these emissions are of the same intrinsic nature. Furthermore, there is no clear evidence for either repeaters or non-repeaters to be preferentially associated with PRSs (Law et al., 2022), indicating that these are two separate aspects of the central engine and its environment (Bhandari et al., 2023).
@@
Apart from their actively repeating behaviour and association with a compact PRS, these three FRB sources also exhibit large values of host galaxy dispersion measure (${\rm DM}\sim 150-450\,{\rm pc\,cm^{-3}}$) and rotation measure (${\rm RM}\sim 10^{3}-10^{5}\,{\rm rad\,m^{-2}}$), indicating the presence of a dense magneto-ionic environment near the PRS emission region. This is consistent with synchrotron radiation originating from a compact magnetized nebula surrounding a young NS, embedded behind an expanding SN ejecta shell. If the observed RM primarily arises from the PRS region, the PRS luminosity should be correlated with the large source RM (Yang et al., 2020). Recently, Bruni et al. (2023) confirmed such a correlation with the detection of the third, less luminous PRS associated with FRB 201124 having ${\rm RM}\sim 10^{3}\,{\rm rad\,m^{-2}}$. They suggest that for lower values of RM, the PRS radio luminosity will likely fall below the detection threshold of current radio telescopes.
The PRSs associated with FRB 121102 and FRB 190520 demonstrate a relatively flat spectrum, $S_{\nu}\propto\nu^{\alpha}$, for $\sim 1-10\,{\rm GHz}$ with a spectral index of $\alpha=-0.27\pm 0.24$ and $-0.41\pm 0.04$, respectively (Marcote et al., 2017; Niu et al., 2022). Both these PRSs have a spectral radio luminosity of $(2-4)\times 10^{29}\,{\rm erg\,s^{-1}\,Hz^{-1}}$ and a flux density $\sim 200\,{\rm\mu Jy}$ at $3\,{\rm GHz}$ (Chatterjee et al., 2017; Niu et al., 2022). In contrast, FRB 201124’s PRS shows an inverted spectrum from $\nu=6-22\,{\rm GHz}$ with $\alpha=1.00\pm 0.43$, flux density $\sim 20\,{\rm\mu Jy}$ at $15\,{\rm GHz}$ and spectral radio luminosity $5\times 10^{27}\,{\rm erg\,s^{-1}\,Hz^{-1}}$ (Bruni et al., 2023). Long-term variation in the radio flux of PRS can be useful to constrain properties of the source. Rhodes et al. (2023) recently found a $\sim 30\%$ reduction in the flux density of FRB 121102’s PRS at 1.3 GHz over a timescale of three years. Similarly, Zhang et al. (2023) reported a $\sim 20\%$ decrease in the flux density of FRB 190520’s PRS at 3 GHz from 2020 to 2021.
@@
Energy injection into the magnetized nebula surrounding the central NS can take place due to either the rotational energy of a young NS that spins down over time (Cordes & Wasserman, 2016; Connor et al., 2016; Lyutikov et al., 2016; Kashiyama & Murase, 2017), or the release of magnetic energy from NS interior due to flares originating close to the magnetar (Lyubarsky, 2014; Kulkarni et al., 2014; Kumar et al., 2017; Margalit & Metzger, 2018; Zhao & Wang, 2021). The injected energy continually drives the expansion of MWN out to the SN ejecta, and the persistent emission is powered by relativistic electrons heated at the termination shock of the magnetar wind. Efficient conversion of NS rotation/magnetic energy to particle energy in the termination shock region is required to explain the observed quasi-steady radio emission. For rapidly rotating young NS, rotational energy is the primary reservoir which powers the wind nebula as is the case with Galactic PWNe (Tanaka & Takahara, 2010). However, for a decades-old magnetar, the interior NS magnetic energy may be more significant in comparison to its rotational energy, as proposed by Margalit & Metzger (2018) to explain the large RM and PRS luminosity of FRB 121102.
@@
The inferred host galaxy DM contribution at $z=0.241$ for FRB 190520 is ${\rm DM}_{\rm host}\simeq 900\,{\rm pc\,cm^{-3}}$ (Niu et al., 2022), which is a factor of five larger than that of typical FRB host galaxies (James et al., 2022). Zhang et al. (2020) used IllustrisTNG simulations to show that the DM contribution from FRB 190520-like host galaxies at $z\approx 0.2$ is ${\rm DM}_{\rm host}\approx 150\pm 100\,{\rm pc\,cm^{-3}}$, which indicates a significant near-source DM contribution, possibly due to an expanding young SN remnant (Piro & Gaensler, 2018; Zhao & Wang, 2021; Katz, 2022). Furthermore, FRB 190520’s DM decreases with time at the rate $-0.09\pm 0.02\,{\rm pc\,cm^{-3}}\,{\rm day}^{-1}$ (Niu et al., 2022), in contrast to the nearly fixed DM of FRB 121102 (Hessels et al., 2019; Oostrum et al., 2020; Li et al., 2021). Such large near-source DM together with its decreasing trend can be explained by an expanding shocked shell of SN remnant (Piro & Gaensler, 2018; Katz, 2022), whereas the large and decreasing RM (e.g., for FRB 121102, Hilmarsson et al. 2021) arises due to the radiative cooling of electrons injected into the magnetized nebula.
@@
In this work, we propose that FRB sources with an associated PRS are powered by young magnetars embedded in a composite of MWN and SN remnant. The observed quasi-steady emission is generated by synchrotron radiation of the MWN, with the dense SN ejecta contributing a large fraction of the near-source DM. This paper is organised as follows. In Section 2, we discuss the physical model for energy injection from the rotation/magnetar-flare-powered wind region into the magnetized nebula. In Section 3, we describe the combined evolution of the MWN and SN ejecta, which is used to compute synchrotron emission from the magnetized nebula. In Section 4, we present the physical constraints needed to explain the detectable radio emission from the PRSs of all three FRBs. Based on the observed radio spectral energy distribution (SED), light curve and time evolution of the near-source DM for these sources, we constrain the magnetar parameters and that of the SN ejecta in Section 5. We also compare the magnetar parameters derived for the rotation- and magnetar-flare-powered models. We summarize our main results and conclude in Section 6.
@@ 2 Physical model
A highly magnetised pulsar (or magnetar) is likely to remain as a compact remnant once a SN explosion takes place. The rotational and/or magnetic energy extracted by the outgoing relativistic wind is injected into the associated nebula which then powers the synchrotron emission from the SN ejecta, detectable at radio frequencies. Figure 1 shows a schematic diagram of a rapidly rotating young magnetar surrounded by a magnetized nebula and SN ejecta. The spindown/magnetic energy powers outflows that are accelerated in the wind zone between the NS light cylinder and the nebula. The energized electrons once injected into the nebula gyrate along the magnetic fields to drive the persistent radio emission that is associated with the FRB. The resultant non-thermal emission is powered by synchrotron radiation in the magnetized nebula and SN ejecta. While most of the RM originates from the magnetic fields in the nebula, bulk of the near-source DM is accumulated in the SN ejecta. FRB persistent emission is observable once the system becomes optically thin to synchrotron self-absorption in the nebula and free-free absorption in the ejecta.
Figure 1: Schematic picture of a magnetar surrounded by magnetized wind nebula and baryonic SN ejecta. Magnetar flares and/or rotationally powered outflows inject particles and magnetic energy into the nebula. Synchrotron radiation, observed as PRS associated with the FRB, is emitted by energetic electrons gyrating within the magnetized nebula. The radio emission is observable once the system becomes optically thin to various processes.
@@ 2.1 Energy source of quasi-steady emission
Murase et al. (2016) considered a young pulsar/magnetar-driven SN as the source of synchrotron emission from wind nebula. Here we extend their model to explain the persistent radio emission from three sources namely FRBs 121102, 190520 and 201124. We first discuss the energy injection mechanism for these sources, considering both rotation- and magnetic-powered scenarios.
The rotation-powered model has been studied extensively in the context of Crab Nebula (e.g., Connor et al. 2016; Cordes & Wasserman 2016; Lyutikov et al. 2016). The intrinsic energy arising from NS rotation is $\mathcal{E}_{\rm rot,i}=0.175M_{\rm ns}R_{\rm ns}^{2}(2\pi/P_{i})^{2}\approx(1.9\times 10^{52}\,{\rm erg})\,P_{i,-3}^{-2}$, where $M_{\rm ns}=1.4\,M_{\odot}$ is the NS mass, $R_{\rm ns}=12\,{\rm km}$ is the NS radius and $P_{i}$ is its initial spin period. Rotation energy is extracted by unipolar induction with spindown luminosity
$L_{\rm sd}\approx(7.2\times 10^{45}\,{\rm erg/s})\,B_{\rm dip,13}^{2}P_{\rm i,-3}^{-4}\left(1+\frac{t}{t_{\rm sd}}\right)^{-2},$ (1)
where $B_{\rm dip}$ is the dipolar magnetic field and $t_{\rm sd}\approx(0.12\,{\rm yr})\,B_{\rm dip,13}^{-2}P_{\rm i,-3}^{2}$ is the characteristic spindown timescale.
Magnetar powered models are promising as evidenced by the recent detection of FRB 200428 from a Galactic magnetar (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). The magnetic energy $\mathcal{E}_{\rm B,int}\approx B_{\rm int}^{2}R_{\rm ns}^{3}/6=(3\times 10^{49}\,{\rm erg})\,B_{\rm int,16}^{2}$ for a given interior magnetic field $B_{\rm int}$, is expected to be more significant in comparison to the rotation energy for a NS with age $t_{\rm age}\gtrsim{\rm few}\,\times 10\,{\rm yr}$. Comparing the minimum energy, $\mathcal{E}_{\rm FRB,min}\approx f_{b}\mathcal{F}_{\rm FRB}\mathcal{R}_{\rm FRB}t_{\rm age}$, needed to explain a repeating FRB with magnetic energy $\mathcal{E}_{\rm B,int}$ provides a lower limit $B_{\rm int}\gtrsim(10^{15}\,{\rm G})\,f_{b}^{1/2}\mathcal{F}_{\rm FRB,40}^{1/2}\mathcal{R}_{\rm FRB,-3}^{1/2}t_{\rm age,9.5}^{1/2}$ (see Kashiyama & Murase 2017). Here $f_{b}$ is the beaming factor, $\mathcal{F}_{\rm FRB}$ is the FRB fluence, $\mathcal{R}_{\rm FRB}$ is the repetition rate for FRB and $t_{\rm age}$ is the NS age.
Magnetic energy injection has been proposed to explain the PRS associated with FRB 121102 along with its high rotation measure. Margalit & Metzger (2018) considered a power-law energy injection rate given by
@@
[…]

### v2

@@ Abstract
Among more than 1000 known fast radio bursts (FRBs), only five sources – FRBs 20121102A, 20190520B, 20201124A, 20240114A and 20190417A – have confirmed associations with persistent radio sources (PRS). The observed quasi-steady emission is consistent with synchrotron radiation from a composite of magnetar wind nebula (MWN) and supernova (SN) ejecta. Using a phenomenological model that incorporates simplified treatments of the nebular dynamics and particle acceleration, we compute the synchrotron flux by solving kinetic equations for energized electrons, accounting for electromagnetic cascades of electron-positron pairs interacting with nebular photons. Within the framework of our model, the rotation-powered scenario requires a young neutron star (NS) with age $t_{\rm age}\approx 20\,{\rm yr}$, dipolar magnetic field $B_{\rm dip}\approx(3-5)\times 10^{12}\,{\rm G}$ and initial spin period $P_{i}\approx 1.5-3\,{\rm ms}$ in an ultra-stripped SN progenitor to account for emissions from FRBs 20121102A and 20190520B. In contrast, FRB 20201124A requires $t_{\rm age}\approx 10\,{\rm yr}$, $B_{\rm dip}\approx 5.5\times 10^{13}\,{\rm G}$ and $P_{i}\approx 10\,{\rm ms}$ in a conventional core-collapse SN progenitor. For the magnetar-flare-powered model, NS aged $t_{\rm age}\approx 25\,/40\,{\rm yr}$ in a USSN progenitor and $t_{\rm age}\approx 12.5\,{\rm yr}$ in a CCSN progenitor explains the observed flux for FRB 20121102A/20190520B and FRB 20201124A, respectively. Finally, we estimate a minimum NS age $t_{\rm age,min}\sim 1-3\,{\rm yr}$ based on the near-source plasma contribution to observed DM, and $t_{\rm age,min}\sim 6.5-10\,{\rm yr}$ from the absence of radio signal attenuation.
@@
radiation mechanisms: general – supernovae: general – stars: magnetars, winds, outflows – transients: fast radio bursts
pagerange: Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebulaeReferencespubyear: 2025
@@ 1 Introduction
Fast radio bursts (FRBs) are energetic millisecond duration pulses of coherent emission, located at cosmological distances, whose physical origin is still debated after a decade since their discovery (Lorimer et al., 2007; Thornton et al., 2013; Spitler et al., 2014; Cordes and Chatterjee, 2019). Many theoretical models have been proposed to explain the nature of their progenitors, both for repeating and one-off events (see Platts et al. 2019, for a recent review). The detection of Galactic FRB 200428 and its association with SGR J1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), suggests that FRBs can originate from magnetars born from the core collapse of massive stars (see e.g., Murase et al. 2016; Kumar et al. 2017; Bhattacharya 2019; Kumar and Bošnjak 2020; Lu et al. 2022).
@@
In addition to FRBs, rapidly rotating pulsars and magnetars have been proposed as central engines of super-luminous supernovae (SNe), stripped-envelope SNe and rapidly-rising optical transients (Metzger et al., 2015; Kashiyama et al., 2016; Hotokezaka et al., 2017; Margalit and Metzger, 2018). Pulsar wind nebulae (PWNe) are known to be efficient particle accelerators and broadband non-thermal emission has been observed from the nebulae of Galactic PWNe (Tanaka and Takahara, 2010, 2013). Murase et al. (2016) proposed quasi-steady synchrotron emission as counterparts of both FRBs and pulsar/magnetar-driven SNe (including SLSNe), suggesting efficient conversion of rotation and/or magnetic energy to particle energies within the nascent MWN of a young NS (e.g., Gaensler and Slane 2006).
Precise localisation of FRBs can provide meaningful insights on their sources, by identifying plausible multi-wavelength counterpart(s) and revealing information about the central engine and its surrounding environment (Michilli et al., 2018). However, until date, only five confirmed FRB-PRS associations have been made in the sample of localized FRBs; namely those of FRBs 20121102A (Chatterjee et al., 2017), 20190520B (Niu et al., 2022), 20201124A (Bruni et al., 2023), 20240114A (Bruni et al., 2024) and 20190417A (Moroianu et al., 2025). Yang et al. (2016) proposed that persistent radio emission from such sources can be produced due to synchrotron-heating process in a self-absorbed magnetar-powered nebula.
@@
PRSs are continuum radio sources that are significantly bright ($L_{\rm PRS}>10^{29}\,{\rm erg\,s^{-1}\,Hz^{-1}}$) and compact ($R_{\rm PRS}<1\,{\rm pc}$), for them to be related to ongoing star formation in their host galaxy (Nimmo et al., 2022; Dong et al., 2024). The polarisation level is found to be significantly different between the FRB 20121102A bursts and its PRS emission (Michilli et al., 2018; Plavin et al., 2022), thereby ruling out the possibility that these emissions are of the same intrinsic nature. Furthermore, there is no clear evidence for either repeaters or non-repeaters to be preferentially associated with PRSs (Law et al., 2022), indicating that these are two separate aspects of the central engine and its environment (Bhandari et al., 2023).
@@
Apart from their actively repeating behaviour and association with a compact PRS, these FRB sources also exhibit large values of host galaxy dispersion measure (${\rm DM}\sim 150-1200\,{\rm pc\,cm^{-3}}$) and rotation measure (${\rm RM}\sim 10^{3}-10^{5}\,{\rm rad\,m^{-2}}$), indicating the presence of a dense magneto-ionic environment near the PRS emission region. If the observed RM primarily arises from the PRS region, the PRS luminosity should be correlated with the large source RM (Yang et al., 2020). Recently, Bruni et al. (2023, 2024) confirmed such a correlation with the detection of less luminous PRSs associated with FRBs 20201124A and 20240114A, both having ${\rm RM}\sim 10^{2}-10^{3}\,{\rm rad\,m^{-2}}$. They suggest that for lower values of RM, the PRS radio luminosity will likely fall below the detection threshold of current radio telescopes.
@@
Energy injection into the MWN surrounding the central NS can occur due to either the rotational energy of a young NS that spins down over time (Cordes and Wasserman, 2016; Connor et al., 2016; Lyutikov et al., 2016; Kashiyama and Murase, 2017), or the release of magnetic energy from NS interior due to flares originating close to the magnetar (Lyubarsky, 2014; Kumar et al., 2017; Beloborodov, 2017; Margalit and Metzger, 2018; Zhao and Wang, 2021). The injected energy continually drives the expansion of MWN out to the SN ejecta, and the persistent emission is powered by relativistic electrons heated at the termination shock of the magnetar wind. Efficient conversion of NS rotation/magnetic energy to particle energy in the termination shock region is required to explain the observed quasi-steady radio emission. For rapidly rotating young NS, rotational energy is the primary reservoir that powers the wind nebula as is the case with Galactic PWNe (Tanaka and Takahara, 2010). However, for a decades-old magnetar, the NS magnetic energy may be more significant in comparison to its rotational energy, as proposed by Beloborodov (2017) to explain the large RM and PRS luminosity of FRB 20121102A.
@@
The inferred host galaxy DM contributions for FRBs 20190520B and 20190417A are ${\rm DM}_{\rm host}\simeq 900-1200\,{\rm pc\,cm^{-3}}$ (Niu et al., 2022), which are almost five times larger than that of typical FRB host galaxies (James et al., 2022). Zhang et al. (2020) used IllustrisTNG simulations to show that the DM contribution from FRB 20190520B-like host galaxies at $z\approx 0.2$ is ${\rm DM}_{\rm host}\approx 150\pm 100\,{\rm pc\,cm^{-3}}$, which indicates a significant near-source DM contribution, possibly due to an expanding young SN remnant (Yang and Zhang, 2017; Piro and Gaensler, 2018; Zhao and Wang, 2021; Katz, 2022). Furthermore, FRB 20190520B’s DM decreases with time at the rate $0.09\pm 0.02\,{\rm pc\,cm^{-3}}\,{\rm day}^{-1}$ (Niu et al., 2022), in contrast to the nearly fixed DM of FRB 20121102A (Hessels et al., 2019; Oostrum et al., 2020; Li et al., 2021). Such large near-source DM together with its decreasing trend can be explained by an expanding shocked shell of SN remnant (Piro and Gaensler, 2018; Katz, 2022), whereas the large and decreasing RM (e.g., for FRB 20121102A, Hilmarsson et al. 2021) arises due to the radiative cooling of electron-positron pairs injected into the magnetized nebula.
@@
In this work, we propose that FRB sources with an associated PRS are powered by young magnetars embedded in a composite of MWN and SN remnant. Observed quasi-steady emission is generated by synchrotron radiation of the MWN, with the dense SN ejecta contributing a large fraction of the near-source DM. This paper is organised as follows. In Section 2, we discuss the physical model for energy injection from the rotation/magnetar-flare-powered wind region into the magnetized nebula. In Section 3, we describe the combined evolution of the MWN and SN ejecta, used to compute synchrotron emission from the magnetized nebula. In Section 4, we present the physical constraints needed to explain the detectable radio emission from the PRSs. Based on the observed radio spectral energy distribution (SED), light curve and time evolution of the near-source DM for these sources, we constrain the magnetar parameters and that of the SN ejecta in Section 5. We also compare the magnetar parameters derived for the rotation- and magnetar-flare-powered models. We discuss our main results with their implications in Section 6 and conclude with a summary in Section 7. Throughout this work, we use the notation $Q=10^{x}Q_{x}$ in CGS units, unless noted otherwise.
@@ 2 Physical model
A highly magnetised pulsar (or magnetar) is likely to remain as a compact remnant once a SN explosion takes place. The rotational and/or magnetic energy extracted by the outgoing relativistic wind is injected into the associated nebula which then powers the synchrotron emission from the SN ejecta, detectable at radio frequencies. Figure 1 shows a schematic diagram of a rapidly rotating young magnetar surrounded by a magnetized nebula and SN ejecta. The spindown/magnetic energy powers outflows that are accelerated in the wind zone between the NS light cylinder and the nebula. The energized electrons and positrons once injected into the nebula gyrate along the magnetic fields to drive the persistent radio emission that is associated with the FRB. The resultant non-thermal emission is powered by synchrotron radiation in the magnetized nebula and SN ejecta. While most of the RM originates from the magnetic fields in the nebula, bulk of the near-source DM is accumulated in the SN ejecta. FRB persistent emission is observable once the system becomes optically thin to synchrotron self-absorption in the nebula and free-free absorption in the ejecta.
Figure 1: Schematic picture of a magnetar surrounded by magnetized wind nebulae and baryonic SN ejecta. Magnetar flares and/or rotationally powered outflows inject particles and magnetic energy into the nebula. Synchrotron radiation, observed as PRS associated with the FRB, is emitted by energetic electron-positron pairs gyrating within the magnetized nebula. The radio emission is observable once the system becomes optically thin to various processes.
@@ 2.1 Energy source of quasi-steady emission
Murase et al. (2016) considered a young pulsar/magnetar-driven SN as the source of synchrotron emission from wind nebulae, and predicted persistent radio emission as counterparts of FRBs earlier than the discovery of FRB 20121102A. Here we extend their model to explain the persistent radio emission from three sources, namely FRBs 20121102A, 20190520B and 20201124A. We first discuss the energy injection mechanism for these sources, considering both rotation- and magnetic-powered scenarios.
The rotation-powered model has been studied extensively in the context of Crab Nebula (e.g., Connor et al. 2016; Cordes and Wasserman 2016; Lyutikov et al. 2016). The intrinsic energy arising from NS rotation is $\mathcal{E}_{\rm rot,i}=0.175M_{\rm ns}R_{\rm ns}^{2}(2\pi/P_{i})^{2}\approx(1.9\times 10^{52}\,{\rm erg})\,P_{i,-3}^{-2}$, where $M_{\rm ns}=1.4\,M_{\odot}$ is the NS mass, $R_{\rm ns}=12\,{\rm km}$ is the NS radius and $P_{i}$ is its initial spin period. The rotation energy can be extracted by unipolar induction. The initial spindown luminosity is given by
$L_{\rm sd,i}\approx\frac{B_{\rm dip}^{2}(2\pi/P_{i})^{4}R_{\rm ns}^{6}}{4c^{3}}(1+{\rm sin}^{2}\chi)$
$\sim(2.4\times 10^{45}\,{\rm erg/s})\,B_{\rm dip,13}^{2}P_{\rm i,-3}^{-4}$ (1)
where $B_{\rm dip}$ is the dipolar magnetic field and $\chi$ is the angle between the axes of rotation and dipole magnetic field (Gruzinov, 2005; Spitkovsky, 2006; Tchekhovskoy et al., 2013). We assume that $B_{\rm dip}$ and ${\rm sin}^{2}\chi=2/3$ are time independent. The corresponding spindown timescale is $t_{\rm sd}\approx(0.25\,{\rm yr})\,B_{\rm dip,13}^{-2}P_{\rm i,-3}^{2}$. The spindown luminosity is almost constant for $t\lesssim t_{\rm sd}$, whereas it evolves as $L_{\rm sd}\approx L_{\rm sd,i}\times(t/t_{\rm sd})^{-2}$ for $t\gtrsim t_{\rm sd}$.
Magnetar powered models are promising as evidenced by the recent detection of FRB 200428 from a Galactic magnetar (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). The magnetic energy $\mathcal{E}_{\rm B,int}\approx B_{\rm int}^{2}R_{\rm ns}^{3}/6=(3\times 10^{49}\,{\rm erg})\,B_{\rm int,16}^{2}$ for a given interior magnetic field $B_{\rm int}$, is expected to be more significant in comparison to the rotation energy for a NS with age $t_{\rm age}\gtrsim{\rm few}\,\times 10\,{\rm yr}$. Comparing the minimum energy, $\mathcal{E}_{\rm FRB,min}\approx f_{b}\mathcal{F}_{\rm FRB}\mathcal{R}_{\rm FRB}t_{\rm age}$, needed to explain a repeating FRB with magnetic energy $\mathcal{E}_{\rm B,int}$ provides a lower limit $B_{\rm int}\gtrsim(10^{15}\,{\rm G})\,f_{b}^{1/2}\mathcal{F}_{\rm FRB,40}^{1/2}\mathcal{R}_{\rm FRB,-3}^{1/2}t_{\rm age,9.5}^{1/2}$ (see Kashiyama and Murase 2017). Here $f_{b}$ is the beaming factor, $\mathcal{F}_{\rm FRB}$ is the FRB fluence, $\mathcal{R}_{\rm FRB}$ is the repetition rate for FRB and $t_{\rm age}$ is the NS age.
Magnetic energy injection has been proposed to explain the PRS associated with FRB 20121102A along with its large rotation measure. Margalit and Metzger (2018) considered a power-law energy injection rate given by
@@
where $t_{\rm inj}\approx 0.6\,{\rm yr}$ is the time when the onset of energy injection occurs and $\alpha\approx 1.3$ is the power-law index. Here $t_{\rm inj}$ is determined by the timescale for magnetic flux to start leaking out of the magnetar core. Beloborodov (2017) proposed that $t_{\rm inj}$ from flares can be comparable to the pair freeze-out timescale $t_{\pm}\sim 0.1\,{\rm yr}$. Although the energy injection index $\alpha$ can depend on the nature of the FRB source, Margalit and Metzger (2018) showed that $\alpha\gtrsim 1$ adequately explains the radio SED of the PRS associated with FRB 20121102A. This implies that the rate of FRB activity slows down over time or the flares become less energetic on average. Although $L_{\rm mag}$ here is modeled with a smooth function of time, the energy release from magnetars can also occur intermittently. For time $t<t_{\rm inj}$, energy injection is dominated by rotation and the total injected energy is $\mathcal{E}_{\rm tot}(t<t_{\rm inj})=\int_{0}^{t}L_{\rm sd}(t)dt$. In contrast, the interior magnetic energy leaks out into the nebula for $t\geq t_{\rm inj}$, and therefore $\mathcal{E}_{\rm tot}(t\geq t_{\rm inj})=\int_{0}^{t}[L_{\rm sd}(t)+L_{\rm mag}(t)]dt$.